# The Sample-Communication Complexity Trade-off in Federated Q-Learning

### v1

@@
Abstract
@@
Keywords: federated Q-learning, communication complexity, sample complexity, trade-offs
@@ 1 Introduction
Reinforcement Learning (RL) (Sutton and Barton, 2018) refers to a paradigm of sequential decision making where an agent aims to learn an optimal policy, i.e., a policy that maximizes the long-term total reward, through repeated interactions with an unknown environment. RL finds applications across a diverse array of fields including, but not limited to, autonomous driving, games, recommendation systems, robotics and Internet of Things (IoT) (Kober et al., 2013; Yurtsever et al., 2020; Silver et al., 2016; Lim et al., 2020).
The primary hurdle in RL applications is often the high-dimensional nature of the decision space that necessitates the learning agent to have to access to an enormous amount of data in order to have any hope of learning the optimal policy. Moreover, the sequential collection of such an enormous amount of data through a single agent is extremely time-consuming and often infeasible in practice (Mnih et al., 2016b). Consequently, practical implementations of RL involve deploying multiple agents to collect data in parallel. This decentralized approach to data collection has fueled the design and development of distributed or federated RL algorithms that can collaboratively learn the optimal policy without actually transferring the collected data to a centralized server, while achieving a linear speedup in terms of the number of agents. Such a federated approach to RL, which does not require the transfer of local data, is gaining interest due to lower bandwidth requirements and lower security and privacy risks. In this work, we focus on federated variants of the vastly popular Q-learning algorithm (Watkins and Dayan, 1992), where the agents collaborate to directly learn the optimal Q-function without forming an estimate of the underlying unknown environment.
A particularly important aspect of designing federated RL algorithms, including federated Q-learning algorithms, is to address the natural tension between sample and communication complexities. At one end of the spectrum lies the naïve approach of running a centralized algorithm with an optimal sample complexity after transferring and combining all the collected data at a central server. Such an approach trivially achieves the optimal sample complexity while suffering from a very high and prohibitive communication complexity. On the other hand, several recently proposed algorithms (Khodadadian et al., 2022; Woo et al., 2023, 2024; Zheng et al., 2024) operate in more practical regimes, offering significantly lower communication complexities when compared to the naïve approach at the cost of sub-optimal sample complexities. These results suggest the existence of an underlying trade-off between sample and communication complexities of federated RL algorithms. The primary goal of this work is to better understand this trade-off in the context of federated Q-learning by investigating these following fundamental questions.
@@ 1.1 Main results
In this work, we consider a setup where $M$ distributed agents — each with access to a local generative model (Kearns and Singh, 1998) — collaborate to learn the optimal Q-function of an infinite-horizon Markov decision process (MDP) (Puterman, 2014), which is defined over a finite state space $\mathcal{S}$ and a finite action space $\mathcal{A}$, and has a discount factor of $\gamma\in(0,1)$. To probe the communication complexity, we consider a common setup in federated learning called the intermittent communication setting (Woodworth et al., 2021), where the agents intermittently share information among themselves with the help of a central server. We provide a complete characterization of the trade-off between sample and communication complexities under the aforementioned setting by providing answers to both the questions. Summarized below, the main result of this work is twofold.
@@
  • Achieving the optimal sample-communication complexity trade-off of federated Q-learning. We propose a new federated Q-learning algorithm called Federated Doubly Variance Reduced Q-Learning (dubbed Fed-DVR-Q), that simultaneously achieves order-optimal sample complexity and communication complexity as dictated by the lower bound. We show that Fed-DVR-Q learns an $\varepsilon$-accurate optimal Q-function in the $\ell_{\infty}$ sense with $\widetilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|}{M\varepsilon^{2}(1-\gamma)^{3}}\right)$ i.i.d. samples from the generative model at each agent while incurring a total communication cost of $\widetilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|}{1-\gamma}\right)$ bits per agent across $\widetilde{\mathcal{O}}\left(\frac{1}{1-\gamma}\right)$ rounds of communication. Thus, Fed-DVR-Q not only improves upon both the sample and communication complexities of existing algorithms, but also is the first algorithm to achieve order-optimal sample and communication complexities (See Table 1 for a comparison).
@@
Q-learning (Li et al., 2024) $1$ $\dfrac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{4}\varepsilon^{2}}$ N/A
Variance Reduced Q-learning (Wainwright, 2019b) $1$ $\dfrac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{3}\varepsilon^{2}}$ N/A
Fed-SynQ (Woo et al., 2023) $M$ $\dfrac{|\mathcal{S}||\mathcal{A}|}{M(1-\gamma)^{5}\varepsilon^{2}}$ $\dfrac{M}{1-\gamma}$
@@
Lower bound ((Azar et al., 2013), This work) $M$ $\dfrac{|\mathcal{S}||\mathcal{A}|}{M(1-\gamma)^{3}\varepsilon^{2}}$ $\dfrac{1}{1-\gamma}$
Table 1: Comparison of sample and communication complexities (per-agent) of various single-agent and federated Q-learning algorithms for learning an $\varepsilon$-optimal Q-function under the synchronous setting. We hide logarithmic factors and burn-in costs for all results for simplicity of presentation. Here, $\mathcal{S}$ and $\mathcal{A}$ represent state and action spaces respectively and $\gamma$ denotes the discount factor. We report the communication complexity only in terms of the number of rounds as existing algorithms assume transmission of real numbers and hence do not report bit-level costs. For the lower bound, Azar et al. (2013) and this work establish the lower bounds for the sample and communication complexities, respectively.
@@ 1.2 Related work
Q-learning has been extensively studied in the single-agent setting in terms of its asymptotic convergence (Jaakkola et al., 1993; Tsitsiklis, 1994; Szepesvári, 1997; Borkar and Meyn, 2000) and its finite-time sample complexity in synchronous (Even-Dar and Mansour, 2004; Kearns and Singh, 1998; Beck and Srikant, 2012; Sidford et al., 2018; Wainwright, 2019a, b; Chen et al., 2020; Li et al., 2024) and asynchronous settings (Chen et al., 2021b; Li et al., 2024, 2021b; Qu and Wierman, 2020; Xia et al., 2024) in terms of convergence in the $\ell_{\infty}$ sense. On the other hand, regret analysis of Q-learning has been carried out in both online settings (Jin et al., 2018; Bai et al., 2019; Ménard et al., 2021; Li et al., 2021a) and offline settings (Shi et al., 2022; Yan et al., 2023), to name a few.
@@
There has also been a considerable effort towards developing distributed and federated RL algorithms. The distributed variants of the classical temporal difference (TD) learning algorithm have been investigated in a series of studies (Chen et al., 2021c; Doan et al., 2019, 2021; Sun et al., 2020; Wai, 2020; Liu and Olshevsky, 2023; Wang et al., 2020; Zeng et al., 2021b). The impact of environmental heterogeneity in federated RL was studied in Wang et al. (2023) for TD learning, and in Jin et al. (2022) when the local environments are known. A distributed version of the actor-critic algorithm was studied by Shen et al. (2023) where the authors established convergence of their algorithm and demonstrated a linear speedup in the number of agents in their sample complexity bound. Chen et al. (2022) proposed a new distributed actor-critic algorithm which improved the dependence of sample complexity on the error $\varepsilon$ with a communication cost of $\widetilde{\mathcal{O}}(\varepsilon^{-1})$Chen et al. (2021a) proposed a communication-efficient distributed policy gradient algorithm with convergence analysis and established a communication complexity of $\mathcal{O}(1/(M\varepsilon))$. Several additional studies (Yang et al., 2023; Zeng et al., 2021a; Lan et al., 2024) have also developed and analyzed other distributed/federated variants of the classical natural policy gradient method (Kakade, 2001)Assran et al. (2019); Espeholt et al. (2018); Mnih et al. (2016a) developed distributed algorithms to train deep RL networks more efficiently.
@@
Federated Q-learning has been explored relatively recently with theoretical sample and communication complexity guarantees. Khodadadian et al. (2022) proposed and analyzed a federated Q-learning algorithm in the asynchronous setting, however, its sample complexity guarantee exhibits pessimistic dependencies with respect to salient problem-dependent parameters. Woo et al. (2023) provided improved analyses for federated Q-learning under both synchronous and asynchronous settings, and introduced importance averaging to tame the heterogeneity of local behavior policies in the asynchronous setting to further improve the sample complexity, showing that a collaborative coverage of the entire state-action space suffices for federated Q-learning. Moving to the offline setting, Woo et al. (2024) proposed a federated Q-learning algorithm for offline RL in the finite-horizon setting and established sample and communication complexities that only require a collaborative coverage of the state-action pairs visited by the optimal policy. Zheng et al. (2024), on the other hand, established a linear speedup for federated Q-learning in the online setting from the regret minimization perspective.
@@
The trade-off between communication complexity and accuracy (or equivalently, sample complexity) has been studied in various federated and distributed learning problems, including stochastic approximation algorithms for convex optimization. Duchi et al. (2014); Braverman et al. (2016) established the celebrated inverse linear relationship between the error and the communication cost for the problem of distributed mean estimation. Similar trade-offs for distributed stochastic optimization, multi-armed bandits and linear bandits have been studied and established across numerous studies, e.g., (Woodworth et al., 2018, 2021; Tsitsiklis and Luo, 1987; Shi and Shen, 2021; Salgia and Zhao, 2023).
@@ 2.1 Markov decision processes
An optimal policy $\pi^{\star}$ is a policy that maximizes the value function uniformly over all the states and it has been shown that such an optimal policy $\pi^{\star}$ always exists (Puterman, 2014). The optimal value and Q-functions are those corresponding to that of an optimal policy $\pi^{\star}$, denoted as $V^{\star}:=V^{\pi^{\star}}$ and $Q^{\star}:=Q^{\pi^{\star}}$ respectively. The optimal Q-function, $Q^{\star}$, is also the unique fixed point of the Bellman optimality operator $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}\times\mathcal{A}$, given by
@@
The popular Q-learning algorithm (Watkins and Dayan, 1992) aims to learn the optimal policy by first learning $Q^{\star}$ as the solution to the above fixed-point equation — via stochastic approximation when $\mathcal{T}$ is only accessed through samples — and then obtaining a deterministic optimal policy via greedy action selection, i.e., $\pi^{\star}(s)=\operatorname*{arg\,max}_{a}Q^{\star}(s,a)$.
@@ 2.2 Performance measures in federated Q-learning
We consider a federated learning setup consisting of $M$ agents, where all the agents face a common, unknown MDP, i.e., the transition kernel and the reward function are the same across agents. In addition, we consider the synchronous setting (Wainwright, 2019a), where each agent has access to an independent generative model or simulator from which they can draw independent samples from the unknown underlying distribution $P(\cdot|s,a)$ for all state-action pair $(s,a)$ (Kearns and Singh, 1998) simultaneously. Let $Z\in\mathcal{S}^{|\mathcal{S}||\mathcal{A}|}$ be a corresponding random vector whose $(s,a)$-th coordinate is drawn from the distribution $P(\cdot|s,a)$, independently of all other coordinates. We define the random operator $\widehat{\mathcal{T}}_{Z}:(\mathcal{S}\times\mathcal{A})\mapsto(\mathcal{S}\times\mathcal{A})$ as
$(\widehat{\mathcal{T}}_{Z}Q)(s,a)=r(s,a)+\gamma V(Z(s,a)),$ (3)
where $V(s^{\prime})=\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$. The operator $\widehat{\mathcal{T}}_{Z}$ can be interpreted as the sample Bellman operator, where we have the relation $\mathcal{T}Q=\mathbb{E}_{Z}\big{[}\widehat{\mathcal{T}}_{Z}Q\big{]}$ for all Q-functions. For a given value of $\varepsilon\in(0,\frac{1}{1-\gamma})$, the objective of agents is to collaboratively learn an $\varepsilon$-optimal estimate (in the $\ell_{\infty}$ sense) of the optimal Q-function of the unknown MDP.

### v2

@@
The Sample-Communication Complexity Trade-off in Federated Q-Learning
Sudeep Salgia
CMU
Department of Electrical and Computer Engineering, Carnegie Mellon University, Pittsburgh, PA 15213, USA.
   Yuejie Chi11footnotemark: 1
CMU
(April 2024)
@@ 1 Introduction
Reinforcement Learning (RL) (Sutton2018RLBook) refers to a paradigm of sequential decision making where an agent aims to learn an optimal policy, i.e., a policy that maximizes the long-term total reward, through repeated interactions with an unknown environment. RL finds applications across a diverse array of fields including, but not limited to, autonomous driving, games, recommendation systems, robotics and Internet of Things (IoT) (Kober2013RoboticsSurvey; Yurtsever2020SelfDriving; Huang2016Go; Lim2020IoT).
The primary hurdle in RL applications is often the high-dimensional nature of the decision space that necessitates the learning agent to have to access to an enormous amount of data in order to have any hope of learning the optimal policy. Moreover, the sequential collection of such an enormous amount of data through a single agent is extremely time-consuming and often infeasible in practice (mnih2016a3c). Consequently, practical implementations of RL involve deploying multiple agents to collect data in parallel. This decentralized approach to data collection has fueled the design and development of distributed or federated RL algorithms that can collaboratively learn the optimal policy without actually transferring the collected data to a centralized server, while achieving a linear speedup in terms of the number of agents. Such a federated approach to RL, which does not require the transfer of local data, is gaining interest due to lower bandwidth requirements and lower security and privacy risks. In this work, we focus on federated variants of the vastly popular Q-learning algorithm (Watkins1992QL), where the agents collaborate to directly learn the optimal Q-function without forming an estimate of the underlying unknown environment.
A particularly important aspect of designing federated RL algorithms, including federated Q-learning algorithms, is to address the natural tension between sample and communication complexities. At one end of the spectrum lies the naïve approach of running a centralized algorithm with an optimal sample complexity after transferring and combining all the collected data at a central server. Such an approach trivially achieves the optimal sample complexity while suffering from a very high and prohibitive communication complexity. On the other hand, several recently proposed algorithms (Khodadadian2022FederatedQL; Woo2023FedSynQ; Woo2024FedOffline; zheng2024federated) operate in more practical regimes, offering significantly lower communication complexities when compared to the naïve approach at the cost of sub-optimal sample complexities. These results suggest the existence of an underlying trade-off between sample and communication complexities of federated RL algorithms. The primary goal of this work is to better understand this trade-off in the context of federated Q-learning by investigating these following fundamental questions.
@@ 1.1 Main results
In this work, we consider a setup where $M$ distributed agents — each with access to a local generative model (Kearns1998FiniteSampleQL) — collaborate to learn the optimal Q-function of an infinite-horizon Markov decision process (MDP) (Puterman2014MDPBook), which is defined over a finite state space $\mathcal{S}$ and a finite action space $\mathcal{A}$, and has a discount factor of $\gamma\in(0,1)$. To probe the communication complexity, we consider a common setup in federated learning called the intermittent communication setting (Woodworth2021IntermittentComm), where the agents intermittently share information among themselves with the help of a central server. We provide a complete characterization of the trade-off between sample and communication complexities under the aforementioned setting by providing answers to both the questions. Summarized below, the main result of this work is twofold.
@@
  • Achieving the optimal sample-communication complexity trade-off of federated Q-learning. We propose a new federated Q-learning algorithm called Federated Doubly Variance Reduced Q-Learning (dubbed Fed-DVR-Q ), that simultaneously achieves order-optimal sample complexity and communication complexity as dictated by the lower bound. We show that Fed-DVR-Q  learns an $\varepsilon$-accurate optimal Q-function in the $\ell_{\infty}$ sense with $\tilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|}{M\varepsilon^{2}(1-\gamma)^{3}}\right)$ i.i.d. samples from the generative model at each agent while incurring a total communication cost of $\tilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|}{1-\gamma}\right)$ bits per agent across $\tilde{\mathcal{O}}\left(\frac{1}{1-\gamma}\right)$ rounds of communication. Thus, Fed-DVR-Q  not only improves upon both the sample and communication complexities of existing algorithms, but also is the first algorithm to achieve order-optimal sample and communication complexities (See Table 1 for a comparison).
@@
Q-learning (Li2023QLMinimax) $1$ $\dfrac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{4}\varepsilon^{2}}$ N/A
Variance Reduced Q-learning (Wainwright2019VarianceReduced) $1$ $\dfrac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{3}\varepsilon^{2}}$ N/A
Fed-SynQ (Woo2023FedSynQ) $M$ $\dfrac{|\mathcal{S}||\mathcal{A}|}{M(1-\gamma)^{5}\varepsilon^{2}}$ $\dfrac{M}{1-\gamma}$
@@
Lower bound ((Azar2013MinimaxPAC), This work) $M$ $\dfrac{|\mathcal{S}||\mathcal{A}|}{M(1-\gamma)^{3}\varepsilon^{2}}$ $\dfrac{1}{1-\gamma}$
Table 1: Comparison of sample and communication complexities (per-agent) of various single-agent and federated Q-learning algorithms for learning an $\varepsilon$-optimal Q-function under the synchronous setting. We hide logarithmic factors and burn-in costs for all results for simplicity of presentation. Here, $\mathcal{S}$ and $\mathcal{A}$ represent state and action spaces respectively and $\gamma$ denotes the discount factor. We report the communication complexity only in terms of the number of rounds as existing algorithms assume transmission of real numbers and hence do not report bit-level costs. For the lower bound, Azar2013MinimaxPAC and this work establish the lower bounds for the sample and communication complexities, respectively.
@@ 1.2 Related work
Q-learning has been extensively studied in the single-agent setting in terms of its asymptotic convergence (Jaakkola1993StochasticIterativeDP; Tsitsiklis1994AsynchronousQL; Szepesvari1997AsymptoticQL; Borkar2000ODE-RL) and its finite-time sample complexity in synchronous (EvenDar2004LearningRates; Kearns1998FiniteSampleQL; Beck2012ConstantStepSize; sidford2018near; Wainwright2019ConeContractive; Wainwright2019VarianceReduced; Chen2020FiniteSampleAnalysis; Li2023QLMinimax) and asynchronous settings (Chen2021Lyapunov; Li2023QLMinimax; Li2021SampleComplexityAsynchronousQL; Qu2020FiniteAsychronousQL; xia2024instance) in terms of convergence in the $\ell_{\infty}$ sense. On the other hand, regret analysis of Q-learning has been carried out in both online settings (jin2018q; bai2019provably; menard2021ucb; li2021breaking) and offline settings (shi2022pessimistic; yan2022efficacy), to name a few.
@@
There has also been a considerable effort towards developing distributed and federated RL algorithms. The distributed variants of the classical temporal difference (TD) learning algorithm have been investigated in a series of studies (Chen2021OffPolicyTDC; Doan2019TD0LinearFunctionMARL; Doan2021DistributedTD0; Sun2020FiniteDecentralizedTDLinear; Wai2020ConvergenceConsensus; liu2023distributed; Wang2020DecentralizedTDTracking; Zeng2021FiniteDecentralizedSA). The impact of environmental heterogeneity in federated RL was studied in Wang2023FederatedTDHeterogeneity for TD learning, and in Jin2022FederatedRLHeterogeneity when the local environments are known. A distributed version of the actor-critic algorithm was studied by Shen2023AsynchronousAdvantageActorCritic where the authors established convergence of their algorithm and demonstrated a linear speedup in the number of agents in their sample complexity bound. Chen2022DecentralizedActorCritic proposed a new distributed actor-critic algorithm which improved the dependence of sample complexity on the error $\varepsilon$ with a communication cost of $\tilde{\mathcal{O}}(\varepsilon^{-1})$Chen2021CommEfficientPolicyGrad proposed a communication-efficient distributed policy gradient algorithm with convergence analysis and established a communication complexity of $\mathcal{O}(1/(M\varepsilon))$.  Xie2023FedKL adopts a distributed policy optimization perspective, which is different from the Q-learning paradigm considered in this work. Moreover, the algorithm in Xie2023FedKL obtains a linear communication cost, which is worse than that obtained in our work. Similarly, Zhang2024FiniteTimeOnPolicy focuses on on-policy learning and incurs a communication cost that depends polynomially on the required error $\varepsilon$. Several additional studies (Yang2023FederatedNPG; Zeng2021DecentralizedNPG; Lan2024AsynchronousPG) have also developed and analyzed other distributed/federated variants of the classical natural policy gradient method (Kakade2001PolicyGradient)Assran2019GossipActorCritic; Espeholt2018Impala; Mnih2016AsynchronousDeepRL developed distributed algorithms to train deep RL networks more efficiently.
@@
Federated Q-learning has been explored relatively recently with theoretical sample and communication complexity guarantees. Khodadadian2022FederatedQL proposed and analyzed a federated Q-learning algorithm in the asynchronous setting, however, its sample complexity guarantee exhibits pessimistic dependencies with respect to salient problem-dependent parameters. Woo2023FedSynQ provided improved analyses for federated Q-learning under both synchronous and asynchronous settings, and introduced importance averaging to tame the heterogeneity of local behavior policies in the asynchronous setting to further improve the sample complexity, showing that a collaborative coverage of the entire state-action space suffices for federated Q-learning. Moving to the offline setting, Woo2024FedOffline proposed a federated Q-learning algorithm for offline RL in the finite-horizon setting and established sample and communication complexities that only require a collaborative coverage of the state-action pairs visited by the optimal policy. zheng2024federated, on the other hand, established a linear speedup for federated Q-learning in the online setting from the regret minimization perspective.
@@
The trade-off between communication complexity and accuracy (or equivalently, sample complexity) has been studied in various federated and distributed learning problems, including stochastic approximation algorithms for convex optimization. Duchi2014DistributedEstimation; Braverman2016CommunicationLowerBounds established the celebrated inverse linear relationship between the error and the communication cost for the problem of distributed mean estimation. Similar trade-offs for distributed stochastic optimization, multi-armed bandits and linear bandits have been studied and established across numerous studies, e.g., (Woodworth2018GraphOracle; Woodworth2021IntermittentComm; Tsitsiklis1987; Shi2021FMAB; Salgia2023LinearBandits).
@@ 2.1 Markov decision processes
An optimal policy $\pi^{\star}$ is a policy that maximizes the value function uniformly over all the states and it has been shown that such an optimal policy $\pi^{\star}$ always exists (Puterman2014MDPBook). The optimal value and Q-functions are those corresponding to that of an optimal policy $\pi^{\star}$, denoted as $V^{\star}:=V^{\pi^{\star}}$ and $Q^{\star}:=Q^{\pi^{\star}}$ respectively. The optimal Q-function, $Q^{\star}$, is also the unique fixed point of the Bellman optimality operator $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}\times\mathcal{A}$, given by
@@
The popular Q-learning algorithm (Watkins1992QL) aims to learn the optimal policy by first learning $Q^{\star}$ as the solution to the above fixed-point equation — via stochastic approximation when $\mathcal{T}$ is only accessed through samples — and then obtaining a deterministic optimal policy via greedy action selection, i.e., $\pi^{\star}(s)=\operatorname*{arg\,max}_{a}Q^{\star}(s,a)$.
@@ 2.2 Performance measures in federated Q-learning
We consider a federated learning setup consisting of $M$ agents, where all the agents face a common, unknown MDP, i.e., the transition kernel and the reward function are the same across agents. In addition, we consider the synchronous setting (Wainwright2019ConeContractive), where each agent has access to an independent generative model or simulator from which they can draw independent samples from the unknown underlying distribution $P(\cdot|s,a)$ for all state-action pair $(s,a)$ (Kearns1998FiniteSampleQL) simultaneously. Let $Z\in\mathcal{S}^{|\mathcal{S}||\mathcal{A}|}$ be a corresponding random vector whose $(s,a)$-th coordinate is drawn from the distribution $P(\cdot|s,a)$, independently of all other coordinates. We define the random operator $\hat{\mathcal{T}}_{Z}:(\mathcal{S}\times\mathcal{A})\mapsto(\mathcal{S}\times\mathcal{A})$ as
$(\hat{\mathcal{T}}_{Z}Q)(s,a)=r(s,a)+\gamma V(Z(s,a)),$ (3)
where $V(s^{\prime})=\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$. The operator $\hat{\mathcal{T}}_{Z}$ can be interpreted as the sample Bellman operator, where we have the relation $\mathcal{T}Q=\mathbb{E}_{Z}\big{[}\hat{\mathcal{T}}_{Z}Q\big{]}$ for all Q-functions. For a given value of $\varepsilon\in(0,\frac{1}{1-\gamma})$, the objective of agents is to collaboratively learn an $\varepsilon$-optimal estimate (in the $\ell_{\infty}$ sense) of the optimal Q-function of the unknown MDP.